the total of 3 cousins' ages is 48. Salley is half as old A Hawkins and 4 years older than Dale. How old are the cousins?
Answer by addingup(3677)

S+H+D = 48
H = 2S
D = S-4
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
S+2S+S-4 = 48
4S-4 = 48
4S = 52
S = 52/4
S = 13
Therefore:
H = 13(2)= 26
D = 13-4 = 9
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
13+26+9 = 48 we have the correct answer.